# Improved Generalization Bounds for Communication Efficient Federated Learning

### v1

@@ 5 Representation Learning Interpretation of R-Round Generalization Bound
The complexity of the model class and the number of samples and local steps used in every round are crucial to minimize the generalization error bound especially in non-iid case in (8) noting that the generalization error bound is loose in comparison to iid setup.
@@ 7 Experimental Results
In this section, we assess the performance of FedALS using ResNet-$20$ as a deep neural network architecture and OPT-$125$M as a large language model. We treat the convolutional layers of ResNet-$20$ as the representation extractor and the final dense layers as the model head. For OPT-$125$M, we consider the first 10 layers of the model as the representation extractor. We used the datasets CIFAR-$10$, CIFAR-$100$, SVHN, and MNIST for image classification and the Multi-Genre Natural Language Inference (MultiNLI) corpus for the LLM. The experimentation was conducted on a network consisting of five nodes alongside a central server. For image classification, we utilized a batch size of $64$ per node. SGD with momentum was employed as the optimizer, with the momentum set to $0.9$, and the weight decay to $10^{-4}$. For the LLM fine-tuning, we employed a batch size of 16 sentences from the corpus, and the optimizer used was AdamW. In all the experiments to perform a grid search for the learning rate, we conducted each experiment by multiplying and dividing the learning rate by powers of two, stopping each experiment after reaching a local optimum learning rate.
@@
The experiments are conducted on Ubuntu $20.04$ using $36$ Intel Core i9-10980XE processors and $5$ GeForce RTX 2080 graphics cards. We repeat each experiment $20$ times and present the error bars associated with the randomness of the optimization. In every figure, we include the average and standard deviation error bars.
@@ 7.2 FedALS in iid Setting
The results for the iid setting are presented in Fig. 2(b), 3(b). In order to obtain these results, the data is shuffled, and then evenly divided among nodes. We note that in this situation, the performance improvement of FedALS is negligible. This is expected since there is a factor of $\frac{1}{K^{2}}$ in the generalization in this case, ensuring that we will have nearly the same population risk as the empirical risk. Therefore, the deciding factor here is the optimization of the empirical risk, which is improved with a smaller $\tau$ as discussed in Remark 5.1. Thus, the improvement of the generalization error using FedALS approach is negligible in this setup.
@@ 7.4 The Role of $\alpha$ and Communication Overhead
As shown in Table 2, it becomes evident that when we increase $\alpha$ from $1$ (FedAvg), we initially witness an enhancement in accuracy owing to improved generalization. However, beyond a certain threshold ($\alpha=10$), further increment in $\alpha$ ceases to contribute to performance improvement. This is due to the adverse impact of a high number of local steps on optimization performance. The trade-off we discussed in the earlier sections is evident in this context. We have also demonstrated the impact of FedALS on the communication overhead in this table.
@@ 7.5 Different Combinations of $\boldsymbol{\phi},\boldsymbol{h}$
In Table 3, we have presented the results of our experiments, illustrating how different combinations of $\boldsymbol{\phi}$ and $\boldsymbol{h}$ influence the model performance in FedALS. The parameter $L$ indicates the number of layers in the model considered as the representation extractor ($\boldsymbol{\phi}$), while the remaining layers are considered as $\boldsymbol{h}$. We observe that for ResNet-$20$, choosing $\boldsymbol{\phi}$ to be the first $16$ layers and performing less aggregation for that seems to be the most effective option.
@@ 8 Conclusion
In this paper, we first characterized generalization error bound for one- and R-round federated learning. One-round generalization bound is tighter than the state of the art. To the best of our knowledge, we designed the R-round generalization bound for heterogeneous data for the first time in the literature. Based on our improved generalization bound analysis and our representation learning interpretation of this analysis, we showed for the first time that less frequent aggregations, hence more local updates, for the representation extractor (usually corresponds to initial layers) leads to the creation of more generalizable models, particularly for non-iid scenarios. This insight led us to develop the FedALS algorithm, which centers around the concept of increasing local steps for the initial layers of the deep learning model while conducting more averaging for the final layers. The experimental results demonstrated the effectiveness of FedALS in a heterogeneous setup.
@@
Putting 10, and 12 together, by the definition of the expected generalization error we get
@@
$\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\frac{L\mathcal{K}(k)^{2}}{\mu}\underbrace{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}_{\text{Expected local generalization}}+2\sqrt{\frac{L}{\mu}}\mathcal{K}(k)\underbrace{\sqrt{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\delta_{k,\mathcal{A}}(\boldsymbol{S})}}_{\text{Root of expected Non-iidness}}\underbrace{\sqrt{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}}_{\text{Root of Expected local generalization}}\bigg{]}$ (14)
@@
We again consider $\boldsymbol{S}_{k}^{\prime}=(\boldsymbol{z}_{{k},{1}}^{\prime},...,\boldsymbol{z}_{{k},{n_{k}}}^{\prime})$, where $\boldsymbol{z}_{{k},{i}}^{\prime}$ is sampled from $\mathcal{D}_{k}$. Let also define $\boldsymbol{S}^{\prime}=(\boldsymbol{S}_{1}^{\prime},...,\boldsymbol{S}_{K}^{\prime})$. Based on Lemma A.1, we can express the expected generalization error as
@@
where .,.,.$\langle.,.\rangle,\|.\|^{2}$ indicate Euclidean inner product, and squared $L2$-norm. Note that (16) holds due to
$f(\boldsymbol{y})\leq f(\boldsymbol{x})+\langle\nabla f(\boldsymbol{x}),\boldsymbol{y}-\boldsymbol{x}\rangle+\frac{L}{2}\|\boldsymbol{y}-\boldsymbol{x}\|^{2}$ (17)
We can bound expectation of the inner product term on the right hand side of (16) using Cauchy–Schwarz inequality as
@@
Now Let’s find an upper bound for $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\mathcal{A}(\boldsymbol{S})-\mathcal{A}(\boldsymbol{S}^{(k)})\|^{2}$ that appears on the right hand side of both (16), and (20). We obtain
@@
$=\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\frac{2\mathcal{K}(k)^{2}}{\mu}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})$ (25)
@@
(23), (24) are based on the definition of local empirical and population risk.
Now we bound $\operatorname{\mathbb{E}}_{k\sim\mathcal{K},\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\nabla\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}l(\mathcal{A}(\boldsymbol{S}^{(k)}),\boldsymbol{z}_{{k},{i}}^{\prime})\|^{2}$ on the right hand side of (20). Note that
@@
where 27 is obtained using the fact that for any $L$-smooth function $f$, we have
@@
and the fact that $\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime})$ is derived from the local ERM, i.e., $\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime})=\arg\min_{M}\sum_{i=1}^{n_{k}}l(M,\boldsymbol{z}_{{k},{i}}^{\prime})$. 28 is based on the definition of local empirical risk.
@@
In (33) we have used the definition of $\delta_{k,\mathcal{A}}(\boldsymbol{S})=\bigg{[}R_{\boldsymbol{S}_{k}}(\mathcal{A}(\boldsymbol{S}))-R_{\boldsymbol{S}_{k}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}))\bigg{]}$. This completes the proof and provide the following upper bound for $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\Delta_{\mathcal{A}}(\boldsymbol{S})$
@@
$+{\sqrt{\frac{8L}{\mu}}}\mathcal{K}(k)\sqrt{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\bigg{(}\delta_{k,\mathcal{A}}(\boldsymbol{S})+\epsilon_{k}(\boldsymbol{S}_{k})\bigg{)}\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\bigg{(}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})+2\epsilon_{k}(\boldsymbol{S}_{k})\bigg{)}}\bigg{]},.$
where $\epsilon_{k}(\boldsymbol{S}_{k})=R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime}))-R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}^{\prime}))$.
@@
First, the new upper bound for $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\mathcal{A}(\boldsymbol{S})-\mathcal{A}(\boldsymbol{S}^{(k)})\|^{2}$ that appears on the right hand side of both (16), and (20). We have
@@
$=\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\frac{4\mathcal{K}(k)^{2}}{\mu}\bigg{(}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})+2R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime}))-2R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}^{\prime}))\bigg{)}$ (45)
@@
Secondly, the bound for $\operatorname{\mathbb{E}}_{k\sim\mathcal{K},\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\nabla\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}l(\mathcal{A}(\boldsymbol{S}^{(k)}),\boldsymbol{z}_{{k},{i}}^{\prime})\|^{2}$ on the right hand side of (20). We get
@@
$\leq 2L\bigg{(}\delta_{k,\mathcal{A}}(\boldsymbol{S}^{(k)})+R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime}))-R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}^{\prime}))\bigg{)},$ (50)
Let’s define $\epsilon_{k}(\boldsymbol{S}_{k})=R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime}))-R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}^{\prime}))$. Putting (16) into (15) and considering (20) we get
@@
Let $l(M_{\boldsymbol{\theta}},\boldsymbol{z})$ be $\mu$-strongly convex and $L$-smooth in $M_{\boldsymbol{\theta}}$. Local models at round $r$ are calculated by doing $\tau$ local steps and the gradient variance is bounded by $\sigma^{2}$. The aggregated model at round $r$ is $M_{\hat{{\boldsymbol{\theta}}}_{r}}$ is obtained by performing FedAvg and where the data points used in round $r$ (i.e., $Z_{k,r}$) are sampled without replacement. The average generalization error, $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\Delta_{FedAvg}(\boldsymbol{S})$, is bounded by
@@
where $A=\tilde{O}\big{(}\sqrt{\frac{\mathcal{C}(M_{\boldsymbol{\theta}})}{|Z_{k,r}|}}+\frac{\sigma^{2}}{\mu\tau}+\frac{L}{\mu}\big{)}$, $B=\tilde{O}\big{(}\operatorname{\mathbb{E}}_{\{Z_{k,r}\}_{k=1}^{K}}\delta_{k,\mathcal{A}}(\{Z_{k,r}\}_{k=1}^{K})+\frac{\sigma^{2}}{\mu\tau}+\frac{L}{\mu}\big{)}$, and $\mathcal{C}(M_{\boldsymbol{\theta}})$ shows the complexity of the model class of $M_{\boldsymbol{\theta}}$.
@@
$\leq\frac{1}{R}\sum_{r=1}^{R}\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{n_{k}}\}}\frac{2L\mathcal{K}(k)^{2}}{\mu}\bigg{(}\Delta_{\mathcal{A}_{k}}(Z_{k,r})+2\epsilon_{k}(Z_{k,r})\bigg{)}$ (56)
$\mathllap{+{\sqrt{\frac{8L}{\mu}}}\mathcal{K}(k)\sqrt{\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\bigg{(}\delta_{k,\mathcal{A}}(\{Z_{k,r}\}_{k=1}^{K})+\epsilon_{k}(Z_{k,r})\bigg{)}\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{n_{k}}\}}\bigg{(}\Delta_{\mathcal{A}_{k}}(Z_{k,r})+\epsilon_{k}(Z_{k,r})\bigg{)}}\bigg{]}}$
@@
$\mathllap{+{\sqrt{\frac{8L}{\mu}}}\mathcal{K}(k)\sqrt{\tilde{O}\bigg{(}\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\delta_{k,\mathcal{A}}(\{Z_{k,r}\}_{k=1}^{K})+\frac{\sigma^{2}}{\mu\tau}+\frac{L}{\mu}\bigg{)}\tilde{O}\bigg{(}\sqrt{\frac{\mathcal{C}(M_{\boldsymbol{\theta}})}{|Z_{k,r}|}}+\frac{\sigma^{2}}{\mu\tau}+\frac{L}{\mu}\bigg{)}}\bigg{]}}$

### v3

@@ 5 Representation Learning Interpretation of R-Round Generalization Bound
The complexity of the model class and the number of samples and local steps used in every round are crucial to minimizing the generalization error bound especially in non-iid case in (9) where the generalization error bound is loose in comparison to iid setup.
@@ 7 Experimental Results
In this section, we assess the performance of FedALS using ResNet-$20$ as a deep neural network architecture and OPT-$125$M as a large language model. We treat the convolutional layers of ResNet-$20$ as the representation extractor and the final dense layers as the model head. For OPT-$125$M, we consider the first 10 layers of the model as the representation extractor. We used the datasets CIFAR-$10$, CIFAR-$100$, SVHN, and MNIST for image classification and the Multi-Genre Natural Language Inference (MultiNLI) corpus for the LLM. The experimentation was conducted on a network consisting of five nodes alongside a central server. For image classification, we utilized a batch size of $64$ per node. SGD with momentum was employed as the optimizer, with the momentum set to $0.9$, and the weight decay to $10^{-4}$. For the LLM fine-tuning, we employed a batch size of 16 sentences from the corpus, and the optimizer used was AdamW. In all the experiments, to perform a grid search for the learning rate, we conducted each experiment by multiplying and dividing the learning rate by powers of two, stopping each experiment after reaching a local optimum learning rate. We repeat each experiment $20$ times and present the error bars associated with the randomness of the optimization. In every figure, we include the average and standard deviation error bars.
@@ 7.2 FedALS in iid Setting
The results for the iid setting are presented in Fig. 2(b), 3(b). In order to obtain these results, the data is shuffled, and then evenly divided among nodes. We note that in this situation, the performance improvement of FedALS is negligible. This is expected since there is a factor of $\frac{1}{K^{2}}$ in the generalization in this case, ensuring that we will have nearly the same population risk as the empirical risk. Therefore, the deciding factor here is the optimization of the empirical risk, which is improved with a smaller $\tau$ as discussed in Remark 5.1. Thus, the improvement of the generalization error using the FedALS approach is negligible in this setup.
@@ 7.4 The Role of $\alpha$ and Communication Overhead
As shown in Table 2, it becomes evident that when we increase $\alpha$ from $1$ (FedAvg), we initially witness an enhancement in accuracy owing to improved generalization. However, beyond a certain threshold ($\alpha=10$), further increment in $\alpha$ ceases to contribute to performance improvement. This is due to the adverse impact of a high number of local steps on optimization performance indicated in Remark 5.1. The trade-off we discussed in the earlier sections is evident in this context. We have also demonstrated the impact of FedALS on the communication overhead in this table.
@@ 7.5 Different Combinations of $\boldsymbol{\phi},\boldsymbol{h}$
In Table 3, we have presented the results of our experiments, illustrating how different combinations of $\boldsymbol{\phi}$ and $\boldsymbol{h}$ influence the model performance in FedALS. The parameter $L$ indicates the number of layers in the model considered as the representation extractor ($\boldsymbol{\phi}$), while the remaining layers are considered as $\boldsymbol{h}$. We observe that for ResNet-$20$, choosing $\boldsymbol{\phi}$ to be the first $16$ layers and performing less aggregation for them seems to be the most effective option.
@@ 8 Conclusion
In this paper, we first characterized generalization error bound for one- and R-round federated learning. One-round generalization bound is tighter than the state of the art. Based on our improved generalization bound analysis and our representation learning interpretation of this analysis, we showed for the first time that less frequent aggregations, hence more local updates, for the representation extractor (usually corresponds to initial layers) leads to the creation of more generalizable models, particularly for non-iid scenarios. This insight led us to develop the FedALS algorithm, which centers around the concept of increasing local steps for the initial layers of the deep learning model while conducting more averaging for the final layers. The experimental results demonstrated the effectiveness of FedALS in heterogeneous setups.
@@
Putting 11, and 13 together, and by the definition of the expected generalization error, we get
@@
$\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\frac{L\mathcal{K}(k)^{2}}{\mu}\underbrace{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}_{\text{Expected local generalization}}+2\sqrt{\frac{L}{\mu}}\mathcal{K}(k)\underbrace{\sqrt{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\delta_{k,\mathcal{A}}(\boldsymbol{S})}}_{\text{Root of expected non-iidness}}\underbrace{\sqrt{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}}_{\text{Root of expected local generalization}}\bigg{]},$ (15)
@@
We again consider $\boldsymbol{S}_{k}^{\prime}=(\boldsymbol{z}_{{k},{1}}^{\prime},...,\boldsymbol{z}_{{k},{n_{k}}}^{\prime})$, where $\boldsymbol{z}_{{k},{i}}^{\prime}$ is sampled from $\mathcal{D}_{k}$. Let also define $\boldsymbol{S}^{(k)}=(\boldsymbol{S}_{1},...,\boldsymbol{S}_{k}^{\prime},...,\boldsymbol{S}_{K})$. Based on Lemma A.1, we can express the expected generalization error as
@@
where $\langle\cdot,\cdot\rangle,\|\cdot\|^{2}$ indicate Euclidean inner product, and squared $L2$-norm. Note that (17) holds due to
$f(\boldsymbol{y})\leq f(\boldsymbol{x})+\langle\nabla f(\boldsymbol{x}),\boldsymbol{y}-\boldsymbol{x}\rangle+\frac{L}{2}\|\boldsymbol{y}-\boldsymbol{x}\|^{2}.$ (18)
We can bound expectation of the inner product term on the right-hand side of (17) using Cauchy–Schwarz inequality as
@@
Now Let’s find an upper bound for $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\mathcal{A}(\boldsymbol{S})-\mathcal{A}(\boldsymbol{S}^{(k)})\|^{2}$ that appears on the right-hand side of both (17), and (21). We obtain
@@
$=\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\frac{2\mathcal{K}(k)^{2}}{\mu}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k}),$ (26)
@@
(24), (25) are based on local empirical and population risk definitions.
Now we bound $\operatorname{\mathbb{E}}_{k\sim\mathcal{K},\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\nabla\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}l(\mathcal{A}(\boldsymbol{S}^{(k)}),\boldsymbol{z}_{{k},{i}}^{\prime})\|^{2}$ on the right-hand side of (21). Note that
@@
where 28 is obtained using the fact that for any $L$-smooth function $f$, we have
@@
and the fact that $\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime})$ is derived from the local ERM, i.e., $\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime})=\arg\min_{M}\sum_{i=1}^{n_{k}}l(M,\boldsymbol{z}_{{k},{i}}^{\prime})$. 29 is based on the definition of local empirical risk.
@@
In (34) we have used the definition of $\delta_{k,\mathcal{A}}(\boldsymbol{S})=\bigg{[}R_{\boldsymbol{S}_{k}}(\mathcal{A}(\boldsymbol{S}))-R_{\boldsymbol{S}_{k}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}))\bigg{]}$. This completes the proof and provides the following upper bound for $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\Delta_{\mathcal{A}}(\boldsymbol{S})$,
@@
$+{\sqrt{\frac{8L}{\mu}}}\mathcal{K}(k)\sqrt{\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\bigg{(}\delta_{k,\mathcal{A}}(\boldsymbol{S})+\epsilon_{k}(\boldsymbol{S}_{k})\bigg{)}\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\bigg{(}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})+2\epsilon_{k}(\boldsymbol{S}_{k})\bigg{)}}\bigg{]},$
where $\epsilon_{k}(\boldsymbol{S}_{k})=R_{\boldsymbol{S}_{k}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}))-R_{\boldsymbol{S}_{k}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}))$.
@@
First, we use a new upper bound for $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\mathcal{A}(\boldsymbol{S})-\mathcal{A}(\boldsymbol{S}^{(k)})\|^{2}$ that appears on the right-hand side of both (17), and (21). We have
@@
$=\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\frac{4\mathcal{K}(k)^{2}}{\mu}\bigg{(}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})+2R_{\boldsymbol{S}_{k}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}))-2R_{\boldsymbol{S}_{k}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}))\bigg{)},$ (46)
@@
Secondly, we derive a new bound for $\operatorname{\mathbb{E}}_{k\sim\mathcal{K},\{\boldsymbol{S}_{k},\boldsymbol{S}_{k}^{\prime}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\|\nabla\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}l(\mathcal{A}(\boldsymbol{S}^{(k)}),\boldsymbol{z}_{{k},{i}}^{\prime})\|^{2}$ on the right hand side of (21). We get
@@
$\leq 2L\bigg{(}\delta_{k,\mathcal{A}}(\boldsymbol{S}^{(k)})+R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}^{\prime}))-R_{\boldsymbol{S}_{k}^{\prime}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}^{\prime}))\bigg{)}.$ (51)
Let’s define $\epsilon_{k}(\boldsymbol{S}_{k})=R_{\boldsymbol{S}_{k}}(\mathcal{A}_{k}(\boldsymbol{S}_{k}))-R_{\boldsymbol{S}_{k}}(\mathcal{A}^{*}(\boldsymbol{S}_{k}))$. Putting (17) into (16) and considering (21) we get
@@
Let $l(M_{\boldsymbol{\theta}},\boldsymbol{z})$ be $\mu$-strongly convex and $L$-smooth in $M_{\boldsymbol{\theta}}$. Local models at round $r$ are calculated by doing $\tau$ local gradient descent steps (5), and the local gradient variance is bounded by $\sigma^{2}$, i.e., $\operatorname{\mathbb{E}}_{\boldsymbol{z}\sim\mathcal{D}_{k}}\|\nabla l(M_{\boldsymbol{\theta}},\boldsymbol{z})-\operatorname{\mathbb{E}}_{\boldsymbol{z}\sim\mathcal{D}_{k}}\nabla l(M_{\boldsymbol{\theta}},\boldsymbol{z})\|^{2}\leq\sigma^{2}$. The aggregated model at round $r$, $M_{\hat{{\boldsymbol{\theta}}}_{r}}$, is obtained by performing FedAvg, and the data points used in round $r$ (i.e., $Z_{k,r}$) are sampled without replacement. The average generalization error bound is
@@
$\leq\frac{1}{R}\sum_{r=1}^{R}\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{|Z_{k,r}|}\}}\frac{2L\mathcal{K}(k)^{2}}{\mu}\bigg{(}\Delta_{\mathcal{A}_{k}}(Z_{k,r})+2\epsilon_{k}(Z_{k,r})\bigg{)}$ (57)
$\mathllap{+{\sqrt{\frac{8L}{\mu}}}\mathcal{K}(k)\sqrt{\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{|Z_{k,r}|}\}_{k=1}^{K}}\bigg{(}\delta_{k,\mathcal{A}}(\{Z_{k,r}\}_{k=1}^{K})+\epsilon_{k}(Z_{k,r})\bigg{)}\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{|Z_{k,r}|}\}}\bigg{(}\Delta_{\mathcal{A}_{k}}(Z_{k,r})+\epsilon_{k}(Z_{k,r})\bigg{)}}\bigg{]}}$
@@
$\mathllap{+{\sqrt{\frac{8L}{\mu}}}\mathcal{K}(k)\sqrt{\tilde{O}\bigg{(}\operatorname{\mathbb{E}}_{\{Z_{k,r}\sim\mathcal{D}_{k}^{|Z_{k,r}|}\}_{k=1}^{K}}\delta_{k,\mathcal{A}}(\{Z_{k,r}\}_{k=1}^{K})+\frac{\sigma^{2}}{\mu\tau}+\frac{L}{\mu}\bigg{)}\tilde{O}\bigg{(}\sqrt{\frac{\mathcal{C}(M_{\boldsymbol{\theta}})}{|Z_{k,r}|}}+\frac{\sigma^{2}}{\mu\tau}+\frac{L}{\mu}\bigg{)}}\bigg{]}},$